A 14 m long truck is driven alongside a 500 m long train with the same velocity of

25 m/s. The front of the truck is 7 m behind the rear of the train. At a certain instant, the

truck stars accelerating at 0,08 m/s2, while the train starts decelerating at 0,4 m/s2. How

long will it take before the rear of the truck passes the front of the train and what distance

would the truck have covered in this time?


1
Expert's answer
2021-08-16T02:36:59-0400
"7+14+500+vt-0.5(0.4)t^2=vt+0.5(0.08)t^2\\t=46.6\ s\\
d=(25)(46.6)+0.5(0.08)(46.6)^2=1250\ m"